what is the direct variation if y=5 when x=10 find y when x=1
Answer by fractalier(6550)

Direct variation is y = kx.
Plug in the first set of data to get k, and then repeat using second set...
5 = 10k
k = 1/2
Then
y = (1/2)(1) = 1/2